For u= i + j - 2k, v= 3i + j + 2k and w= 2i - 2j - 3k
Will someone please show me how to find the cross product of u and w?
Thanks,
-Nick

Answer by Alan3354(69443)

For u= i + j - 2k, v= 3i + j + 2k and w= 2i - 2j - 3k
Will someone please show me how to find the cross product of u and w?
---------
Use determinants.
|i +j +k|
|1 +1 -2| = cross product
|2 -2 -3|
----------
i(-3 - 4) -j(-3 + 4) + k(-2 - 2)
= -7i -j -4k